# NeuralFLoC: Neural Flow-Based Joint Registration and Clustering of Functional Data

### v1

@@ 1.1 Motivation
The second strategy seeks to jointly perform registration and clustering within a unified statistical framework [Liu and Yang, 2009; Wu and Hitchcock, 2016; Zeng et al., 2019]. Despite their conceptual appeal, existing joint models exhibit notable limitations: (1) they often impose restrictive assumptions on the warping functions, such as simple time shifts [Liu and Yang, 2009], Dirichlet process priors [Wu and Hitchcock, 2016], or parametric mixed-effects structures [Zeng et al., 2019]; (2) they are difficult to extend to multivariate or vector-valued functional data; (3) they rely on computationally expensive inference procedures, limiting scalability to large datasets.
@@ 3.1 Latent Initialization via 1D-CNN
1D convolutional neural networks (1D-CNNs) effectively capture local morphological patterns in functional data [Kiranyaz et al., 2015; Jiang et al., 2025]. Our encoder, parameterized by $\Theta_{\text{enc}}$, defines a nonlinear mapping $\Phi:\mathcal{X}\to\mathcal{Z}$ that projects a raw sequence $\mathbf{x}_{i}\in\mathbb{R}^{1\times T}$ into a latent space $\mathcal{Z}\in\mathbb{R}^{C}$, where $C$ denotes the number of clusters (set via validation or overparameterization with sparsity [Arbelaitz et al., 2013; Peng et al., 2016]). The latent vector $\mathbf{z}_{i}$ is computed as:
@@ 3.2 Continuous Time-Warping via Neural ODEs
$\frac{d\tau_{i}(t)}{dt}=f\!\left(\tau_{i}(t),t,\mathbf{z}_{i};\Theta_{\text{ode}}\right),\quad\tau_{i}(0)=\mathbf{z}_{i},$ (5)
where $f$ is a multi-layer perceptron. A Softplus activation is applied to the output of $f$ to enforce strict monotonicity ($\dot{\gamma}>0$), as required for valid warping [Srivastava and Klassen, 2016]. The normalized warping function is obtained via boundary scaling:
$\hat{\gamma}_{i}(t)=\frac{\tau_{i}(t)-\tau_{i}(0)}{\tau_{i}(1)-\tau_{i}(0)}.$ (6)
@@
Initial assignments $\{\mathbf{p}_{i}\}$ are set uniformly. The aligned function $\tilde{x}_{i}(t)=x_{i}(\gamma_{i}(t))$ is computed using 1D linear interpolation, which ensures stable gradient flow during backpropagation. The complete procedure is outlined in Algorithm 1.
@@ 3.5 Optimization, Complexity, and Extension
10:Output: Optimal warping functions $\{\gamma_{i}\}$ and assignments $P$.
@@ 5.1 Experimental Settings
For a fair comparison, all sequential registration-then-clustering pipelines are evaluated by combining SrvfRegNet with each downstream clustering method (FPCA, B-spline, FAE, RandomNet, K-graph). The variant of our method that uses SrvfRegNet instead of the Neural ODE warper is denoted as Ours (N-ODE $\rightarrow$ SrvfRegNet).
@@ 5.2 Experimental Results
Table 1 summarizes the registration and clustering results across all datasets. In contrast to sequential pipelines (e.g., SrvfRegNet followed by clustering) that risk “over-warping” and do not use cluster structure to inform alignment, our joint framework (Ours) yields superior registration and higher clustering accuracy. Only our full model and its discrete-warping variant (Ours (N-ODE $\to$ SrvfRegNet)) optimize registration and clustering simultaneously. Quantitatively, the complete model outperforms the discrete variant on nearly every dataset in both tasks. Qualitatively, Figure 2(c,d) on the Shapes dataset shows that our method produces smoother warpings and clearer cluster separation, visually justifying the improved metrics. These findings highlight the benefit of continuous-flow warping in the joint framework.
@@
The bottom rows of Table 1 ablate our method’s core components: Neural ODE-based warping with SRVF transformation, and Fourier spectral clustering with soft assignment. Replacing the ODE warper with discrete SrvfRegNet (Ours (N-ODE $\rightarrow$ SrvfRegNet)) causes a consistent performance drop, confirming the advantage of continuous-flow registration. Removing registration entirely (Ours w/o Reg) severely degrades clustering, while disabling clustering (Ours w/o Clu) impairs alignment. As shown in Table A.2 in Appendix C, paired $t$-tests over 10 random seeds confirm that ablating any component yields a statistically significant decline ($p<0.01$ for most comparisons). These results demonstrate that both registration and clustering modules are essential to our framework’s performance.
@@ 6 Discussion
We simulate irregular sampling by applying temporal perturbations to all datasets. First, each raw curve is fitted with a Fourier spline [Wahba, 1990]; then, an uneven grid is generated using the same perturbation method as in the Irregular Sampling scenario of [Zeng and Jiang, 2025]. The resulting evaluations on the perturbed grid are treated as irregularly sampled data. Three perturbation levels are tested ($\sigma_{T}\in\{0.1,0.2,0.3\}$). Without hyperparameter tuning, our method maintains stable registration and clustering performance across all datasets except Shapes, where performance degrades moderately as $\sigma_{T}$ increases (middle panel, Figure 4).

### v2

@@ 1.1 Motivation
The second strategy seeks to jointly perform registration and clustering within a unified statistical framework [Liu and Yang, 2009; Wu and Hitchcock, 2016; Zeng et al., 2019; Gaffney and Smyth, 2004]. Despite their conceptual appeal, existing joint models exhibit notable limitations: (1) they often impose restrictive assumptions on the warping functions, such as simple time shifts [Liu and Yang, 2009], Dirichlet process priors [Wu and Hitchcock, 2016], or parametric mixed-effects structures [Zeng et al., 2019]; (2) they are difficult to extend to multivariate or vector-valued functional data; (3) they rely on computationally expensive inference procedures, limiting scalability to large datasets.
@@ 3.1 Latent Initialization via 1D-CNN
1D convolutional neural networks (1D-CNNs) effectively capture local morphological patterns in functional data [Kiranyaz et al., 2015; Jiang et al., 2025]. Our encoder, parameterized by $\Theta_{\text{enc}}$, defines a nonlinear mapping $\Phi:\mathcal{X}\to\mathcal{Z}$ that projects a raw sequence $\mathbf{x}_{i}\in\mathbb{R}^{1\times T}$ into a latent space $\mathcal{Z}\in\mathbb{R}^{C}$, where $C$ denotes the number of clusters (see section 6 on how to determine $C$). The latent vector $\mathbf{z}_{i}$ is computed as:
@@ 3.2 Continuous Time-Warping via Neural ODEs
$\frac{d\tau_{ij}(t)}{dt}=f\!\left(\tau_{ij}(t),t,{z}_{ij};\Theta_{\text{ode}}\right),\quad\tau_{ij}(0)={z}_{ij},$ (5)
where $f$ is a multi-layer perceptron, and $z_{ij}$ is the $j$-th element of $\bm{z}_{i}$. A Softplus activation is applied to the output of $f$ to enforce strict monotonicity ($\dot{\gamma}>0$), as required for valid warping [Srivastava and Klassen, 2016]. The normalized warping function is obtained via boundary scaling:
$\hat{\gamma}_{ij}(t)=\frac{\tau_{ij}(t)-\tau_{ij}(0)}{\tau_{ij}(1)-\tau_{ij}(0)}.$ (6)
@@
Initial assignments $\{\mathbf{p}_{i}\}=\{p_{i1},...,p_{iC}\}$ are set uniformly. Note that the vector $\bm{z}_{i}\in R^{C}$(Eq. 4) is not a free latent embedding: its $j$-th coordinate initializes the $j$-th cluster-specific ODE flow whose mixture forms the final warp (Eq. 7). Thus, the dimensionality is structural rather than a representational bottleneck: the 1D-CNN encoder itself has full expressive power, while the final linear layer simply reorganizes its outputs into the flow-related C-dimensional structure.The aligned function $\tilde{x}_{i}(t)=x_{i}(\gamma_{i}(t))$ is computed using 1D linear interpolation, which ensures stable gradient flow during backpropagation. The complete procedure is outlined in Algorithm 1.
@@ 3.5 Optimization, Complexity, and Extension
10:Output: Optimal warping functions $\{\gamma_{i}\}$ and assignments $\mathbf{p}_{i}$.
@@ 5.1 Experimental Settings
  • Joint Methods: JPCCA[Gaffney and Smyth, 2004], BSRC[Wu and Hitchcock, 2016] and SRC[Zeng et al., 2019].These are classical staistical methods that perform simultaneous functional registration and clustering.
For a fair comparison, all sequential registration-then-clustering pipelines are evaluated by combining SrvfRegNet with each downstream clustering method (FPCA, B-spline, FAE, RandomNet, K-graph). The variant of our method that uses SrvfRegNet instead of the Neural ODE warper is denoted as Ours (N-ODE $\rightarrow$ SrvfRegNet), and the variant that uses short-time Fourier transform (STFT) features instead of the Fourier basis is denoted as Ours (Fourier $\rightarrow$ STFT).
@@ 5.2 Experimental Results
Table 1 summarizes the registration and clustering results across all datasets. In contrast to sequential pipelines (e.g., SrvfRegNet followed by clustering) that risk “over-warping” and do not use cluster structure to inform alignment, our joint framework (Ours) yields superior registration and higher clustering accuracy. To verify whether these benefits originate from our continuous-flow formulation or merely from joint optimization, we compare our full Neural ODE model with three types of alternatives: a discrete-warping variant that substitutes the ODE with SrvfRegNet but retains joint training (Ours (N-ODE $\to$ SrvfRegNet)); several representative joint registration–clustering methods (BSRC, JPCCA, SRC); and a feature-engineering variant that uses short-time Fourier transform features instead of the Fourier basis (Ours (Fourier $\to$ STFT)). Quantitatively, the complete model outperforms the discrete variant on every dataset, with the most pronounced gap on Wave (d=1). On this dataset, which exhibits minimal phase variability, the discrete variant severely underperforms because jointly training SrvfRegNet without the ODE’s smoothness constraints causes optimization instability, weaker regularization that encourages over-warping even when warps are nearly trivial, and difficulty in adapting pretrained alignment alongside clustering. Our continuous Neural ODE warper avoids these issues entirely, achieving the highest accuracy, and this case therefore clearly highlights the superiority of continuous flows in maintaining stable alignment even when the true warping structure is simple. The classical joint baselines, despite also solving registration and clustering simultaneously, all yield markedly lower clustering accuracy. This demonstrates that the observed gains do not arise from the joint paradigm itself, but from our continuous diffeomorphic warping, mixture-of-flows formulation, and end-to-end optimization. Furthermore, replacing the Fourier basis with STFT features consistently degrades performance. After SRVF-based phase alignment, most residual amplitude variation resides in a low-frequency subspace; STFT overemphasizes high-frequency artifacts from imperfect warping, whereas the truncated Fourier basis provides a natural geometry-consistent denoising effect. Qualitatively, Figure 2(c,d) on the Shapes dataset shows that our method produces smoother warpings and clearer cluster separation, visually justifying the improved metrics. These findings highlight the benefit of continuous-flow warping in the joint framework.
@@
[…]
@@ 6 Discussion
Handling Unknown Cluster Number in Practice.
Although for fair comparison we used the ground-truth cluster number $C$ across all baselines, our method can naturally select $C$ in practice via an elbow analysis on the total variation (TV) metric from Fisher–Rao alignment [Chen and Srivastava, 2021]. The rationale is that alignment quality, as measured by TV, is optimized when the cluster number is correctly chosen. Importantly, TV does not rely on ground-truth labels, making it a principled, method-specific criterion for cluster selection. Applying this elbow analysis to our five datasets yields cluster number estimates of $(3,2,2,2,3)$, closely matching the true cluster numbers $(2,2,2,2,3)$. This demonstrates that our framework offers a data-driven approach for choosing $C$ when it is unknown.
@@
We simulate irregular sampling by applying temporal perturbations to all datasets. First, each raw curve is fitted with a Fourier spline [Wahba, 1990]; then, an uneven grid is generated using the same perturbation method as in the Irregular Sampling scenario of [Zeng and Jiang, 2025]. The resulting evaluations on the perturbed grid are treated as irregularly sampled data. Three perturbation levels are tested ($\sigma_{T}\in\{0.1,0.2,0.3\}$).We apply a standard FDA preprocessing step: project each curve onto a truncated Fourier basis and reconstruct on a common grid. This preserves low-frequency structure—the most relevant component after SRVF alignment—while ensuring consistent input resolution. Without hyperparameter tuning, our method maintains stable registration and clustering performance across all datasets except Shapes, where performance degrades moderately as $\sigma_{T}$ increases (middle panel, Figure 4).